40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 7 03:14:06 2013, Jenne, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
    Reply  Wed Aug 21 23:50:40 2013, Koji, Update, SUS, PRM SUS_LSC violin (FM5) set to correct frequency 
       Reply  Fri Aug 23 01:52:34 2013, Jenne, Update, SUS, Violin filters moved to LSC, from SUS MICHloopNotches_23Aug2013.pdf
          Reply  Fri Aug 23 14:19:28 2013, rana, Update, SUS, Violin filters moved to LSC, from SUS 
Message ID: 9048     Entry time: Wed Aug 21 23:50:40 2013     In reply to: 8533     Reply to this: 9057
Author: Koji 
Type: Update 
Category: SUS 
Subject: PRM SUS_LSC violin (FM5) set to correct frequency 

[Jenne Koji]

It seems that the PRM violin mode freqs shifted from 625-ish to 640Hz.
The peaks rang up because of the servo.

Once the notch freq was shifted to 640Hz, the violin mode started to decay.

ellip("BandStop",4,1,90,636,644) gain(1.12202)

ELOG V3.1.3-